OpenStudy (anonymous):

find the area A and circumference C or a circle of radius 5 meters?

OpenStudy (bacon):

do you know the radius of the circle?

OpenStudy (anonymous):

5 meters

Parth (parthkohli):

\[ Circumference = 2\pi r\\ Area = \pi r^2\]

OpenStudy (anonymous):

78.5

OpenStudy (anonymous):

3.14x25

OpenStudy (anonymous):

right?

Parth (parthkohli):

\[2\times \pi \times 5\implies 10 \times \pi \\ \pi \times 5 \times 5 \implies 25 \times \pi \]

Parth (parthkohli):

Just keep it like \(10\pi \text{ meters}\) and \(25\pi \text{ sq. meters}\)

Parth (parthkohli):

You may multiply by 3.14, yes.

OpenStudy (anonymous):

so the answer is 78.5

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[A=5^2*\pi=25\pi=25*3.14\]the answer is your